Mechanics

A body is just being revolved in a vertical circle of radius R such that lowest point of trajectory just toches the ground. The string breaks when the body is at the highest point. The horizontal distance covered by the body after the string breaks is :-
Options
1. 2R
2. R
3. R(2)^1/2
4. 4R

Sayeda Mahfuja Rahman
8 Years agoGrade

1 Answer

Khimraj
ApprovedApproved Tutor Answer8 Years ago
at highest point velocity will be  v = \sqrt{gR} and it will be parallel to ground.
So time taken to reach ground
2R = (½)gt2
t = 2\sqrt{R/g}
horizontal distance covered = v*t = 2R.
Hope it clears. If you like anser then please approve it.